# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v2

@@ I.5 Beyond neuroscience: diverse applications of the FitzHugh-Nagumo model
Cardiology: In 1962, based on available experimental data, Noble adapted Hodgkin and Huxley’s model of excitable neurons to describe the dynamics of Purkinje fibers in the heart Noble (1962). The key modification to the HH model involved considering potassium current flowing through two nonlinear resistances. Since then, a plethora of cardiac models have been developed, ranging from subcellular to whole organ descriptions, with varying levels of detail.
Among these models, the FHN model stands out as one of the simplest yet widely studied models for capturing general dynamical features of cardiac cells Keener (1988); Clayton and Panfilov (2008); Keener and Sneyd (2009); Van Capelle and Durrer (1980); Franzone et al. (2006); Davidenko et al. (1995); Courtemanche et al. (1990); Keener and Sneyd (2009); Aliev and Panfilov (1996); Biktashev (2002); Pertsov et al. (1993); Nash and Panfilov (2004); Watanabe et al. (2004); Sermesant et al. (2006); Cherubini et al. (2008); Glass and Josephson (1995); Le and Sotiropoulos (2013); Panfilov and Keener (1995); Nagai et al. (2000). Another model in common use, the Aliev-Panfilov model Aliev and Panfilov (1996), adapts the FHN framework to capture the rate-dependent variations in action potential duration.
@@ II.2 Linear stability of the stationary solutions
The position of these bifurcations vary as a function of the control parameters of the system $a,b$ and $\epsilon$, leading to the emergence of different dynamical behaviors and regimes. An example of the distribution of these regimes and the bifurcations which define them is illustrated in the $(a,b)$-phase diagram of Fig. 5B for $\epsilon=0.01$.
@@ III The diffusively coupled FitzHugh-Nagumo system
Originally conceived for neuroscience applications, the FHN model FitzHugh (1961) and its spatial extension Nagumo et al. (1962) have been used beyond their initial scope, finding relevance across various disciplines. In cardiology, the model is celebrated for its ability to simulate spiral wave dynamics Franzone et al. (2006); Davidenko et al. (1995); Courtemanche et al. (1990); Keener (1988); Keener and Sneyd (2009); Aliev and Panfilov (1996); Biktashev (2002); Pertsov et al. (1993); Nash and Panfilov (2004); Watanabe et al. (2004); Sermesant et al. (2006); Cherubini et al. (2008); Glass and Josephson (1995); Le and Sotiropoulos (2013); Panfilov and Keener (1995); Nagai et al. (2000), while more generally in biology, it is recognized for generating traveling waves that effectively convey information Gelens et al. (2014); Noguchi (2023); Nolet et al. (2020a); Gambino et al. (2022, 2023); Cartwright et al. (1999); Duffy et al. (1998); Sneyd et al. (2017b); Lee (1997); Gelens et al. (2015). The model’s utility extends to chemistry and computational sciences Moayeri et al. (2020); Dehghan and Narimani (2018); Fuselier and Wright (2013); Rogers and McCulloch (1994); Kuramoto et al. (1998); Lin et al. (2004), where it aids in understanding complex systems. Beyond its interdisciplinary applications, the FHN model has been instrumental in investigating mathematical properties Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Keener (1986a); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Ohta et al. (1989); Iqbal et al. (2017); Elphick et al. (1990); Meron (1992); Petrich and Goldstein (1994); Hagberg and Meron (1994a); Goldstein et al. (1996); Vasiev (2004); Nekhamkina and Sheintuch (2006); Panfilov and Hogeweg (1993); Winfree (1991a); Zheng and Shen (2015); Tsyganov et al. (2014); Rombouts and Gelens (2020); Winfree (1991b); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Zaritski et al. (2004); Bär and Eiswirth (1993); Kuznetsov et al. (2017); Shepelev et al. (2017); Browne et al. (2008); Guckenheimer and Kuehn (2010, 2009); Perc (2005b); Shima and Kuramoto (2004); Hakim and Karma (1999); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Métens et al. (1997); Zemskov and Epstein (2010); Hastings (1982); Lin et al. (2004); Bode (1997); Tobias and Knobloch (1998); Marée and Panfilov (1997); Shoji et al. (2007); Yang et al. (2003), establishing it as a fundamental model in theoretical studies. This section aims to provide a theoretical exploration into the existence and linear stability of stationary homogeneous solutions and the emergence of spatially structured dynamics such as traveling waves and extended patterns. For in-depth mathematical treatments, readers are directed to Cross and Greenside (2009).
@@ III.6.2 Stationary patterns in two spatial dimensions
Stationary patterns, stemming from Turing bifurcations, have been a significant area of study. Various researchers have looked into the genesis of these spatial formations, formulating amplitude equations Iqbal et al. (2017); Zheng and Shen (2015). A notable challenge is the close proximity of Turing and Hopf bifurcations, complicating pattern identification. To circumvent this, strategies involve operating near the Turing threshold or employing white noise to steer the system towards stable pattern branches. Our numerical simulations reveal a diversity of stable spatial patterns across different control parameters, specifically for $b=1.26$, $\epsilon=0.5$, $D_{u}=1$, and $D_{v}=5$, as illustrated in Fig. 13. These patterns, transitioning from negative to positive spots through labyrinthine structures, underscore the labyrinth configuration’s robust stability for the chosen parameters. Notably, Goldstein’s work on interface growth within the FHN framework Goldstein et al. (1996), inspired by chemical front interactions Lee et al. (1993), aligns with our findings in the labyrinth domains (Fig. 13B).
@@ III.6.5 Spiral waves
For the latter, we focus on spirals, which have been widely studied in the FHN model for mathematical purposes Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) and for their applicability in cardiology Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), where spiral behavior has been associated with arrhythmias. From the simple spirals studied by Erhaud and Winfree Meron (1992); Winfree (1991b) to the complex collisions studied by the latter Winfree (1991a), and through spiral turbulence emerging from the loss of stability of these structures Zaritski et al. (2004); Bär and Eiswirth (1993), the FHN model has been utilized to characterize several regimes associated with spiral behavior. In Fig. 14B, we illustrate the creation of a stable spiral arising from a phase defect induced by considering horizontal gradient-like initial conditions for $u$ and horizontal gradient-like initial conditions for $v$ in an oscillatory regime ($a=0.1$, $b=1$, $\varepsilon=0.005$, $D_{u}=1$, $D_{v}=1$).
Spiral wave patterns, particularly relevant in cardiological contexts due to their association with arrhythmic behaviors Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), represent another facet of the FHN model’s versatility. From the foundational works on simple spirals Meron (1992); Winfree (1991b) to the studies on spiral turbulence Zaritski et al. (2004); Bär and Eiswirth (1993) and collision dynamics Winfree (1991a), the model serves as a vital tool in understanding the nuances of spiral behavior. Fig. 14B demonstrates the formation of a stable spiral, a phenomenon induced by specific initial conditions (here a phase defect) within an oscillatory setting. This insight into spiral formation within the FHN model not only advances mathematical explorations Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) but also has profound implications in the study of cardiac arrhythmias Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), offering potential avenues for therapeutic interventions.
@@ IV Discretely coupled FitzHugh-Nagumo equations
F Goetze and Lai (2022); Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Stefanescu and Jirsa (2008); Li et al. (2021); Binczak et al. (2006); Yanagita et al. (2005) F T Wang et al. (2009); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Toral et al. (2003b); Yu et al. (2021); Brandstetter et al. (2010); Panchuk et al. (2013); Ciszak et al. (2003); Schöll et al. (2009); Saha and Feudel (2017) T F Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner (2004); Brown et al. (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005a); Takembo et al. (2019) F Lei and Yang (2021); Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007); Steur et al. (2009); Zhang and Liao (2017); Chouzouris et al. (2018); Zaks et al. (2005); Vodeneev et al. (2011); Sukhov et al. (2011); Schmidt et al. (2017); García-Ojalvo and Schimansky-Geier (1999); Cartwright (2000); Gerster et al. (2020); Scialla et al. (2021) T Burić and Todorović (2003); Kori et al. (2012); Perlikowski et al. (2010) T Sawicki et al. (2019); Schülen et al. (2019) Discretely coupled ODEs, describing various network configurations, have garnered substantial attention in diverse fields such as neuroscience Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020), electrical systems Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006); Takembo et al. (2019); Ciszak et al. (2003), and biology Kori et al. (2012); Vodeneev et al. (2011); Sukhov et al. (2011); Cartwright (2000), owing to their relevance and the broad spectrum of topologies and coupling terms they accommodate. This versatility allows for the exploration of various phenomena, enriching the mathematical study of networks Vaidyanathan (2015); Hoff et al. (2014); Lei and Yang (2021); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Lindner (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005a); Steur et al. (2009); Zaks et al. (2005); Schmidt et al. (2017); García-Ojalvo and Schimansky-Geier (1999); Gerster et al. (2020); Yanagita et al. (2005).
@@ V Conclusions
We structured our analysis into three primary sections. Firstly, we examined the original FHN model (Eq. 8), discussing widely observed dynamical regimes such as monostability, multistability, relaxation oscillations, and excitability. We examined the role of local and global bifurcations in shaping these regimes, emphasizing the importance of time scale separation. Secondly, we explored the diffusively coupled FHN model (Eq. 10), introducing spatial coupling through diffusion. Through theoretical analysis, we investigated stationary homogeneous solutions, their linear stability, and spatially structured dynamic solutions, including traveling structures and spatially extended patterns. We studied the emergence of a Turing instability and the resulting spatially structured Turing patterns. Additionally, we examined front solutions, localized states, traveling pulses, and pacemaker-driven waves within the oscillatory domain, highlighting the richness of patterns that arose in different spatial dimensions. Lastly, we explored discretely coupled FHN equations (Eq. 11). This is the broadest category as here one can consider a multitude of different network topologies and coupling terms. We focussed ourselves on synchronization properties in two coupled FHN modules, the existence of traveling waves when transitioning from continuous diffusive coupling to discrete coupling, and the emergence of chimera states characterized by spatiotemporal patterns of coherent and incoherent behavior.

### v3

@@ I.5 Beyond neuroscience: diverse applications of the FitzHugh-Nagumo model
Cardiology: In 1962, based on available experimental data, Noble adapted Hodgkin and Huxley’s model of excitable neurons to describe the dynamics of Purkinje fibers in the heart Noble (1962). The key modification to the HH model involved considering potassium current flowing through two nonlinear resistances. Since then, a plethora of cardiac models have been developed, ranging from subcellular to whole organ descriptions, with varying levels of detail Alonso et al. (2016).
Among the various models used to study cardiac cells, the FHN model stands out as one of the simplest yet most widely studied for capturing the general dynamical features of cardiac cells Keener (1988); Clayton and Panfilov (2008); Keener and Sneyd (2009); Van Capelle and Durrer (1980); Franzone et al. (2006); Davidenko et al. (1995); Courtemanche et al. (1990); Keener and Sneyd (2009); Aliev and Panfilov (1996); Biktashev (2002); Pertsov et al. (1993); Nash and Panfilov (2004); Watanabe et al. (2004); Sermesant et al. (2006); Cherubini et al. (2008); Glass and Josephson (1995); Le and Sotiropoulos (2013); Panfilov and Keener (1995); Nagai et al. (2000); Alonso et al. (2016); Bär and Eiswirth (1993b); Alonso et al. (2003). In addition to the FHN model, several other generic models are used to reproduce the qualitative properties of cardiac excitation. The Barkley model, which is very similar to the FHN model, is widely used in numerical studies due to its efficient numerical implementation Dowle et al. (1997); Barkley (1991). The Bär–Eiswirth model extends the Barkley model and was introduced specifically to study spiral breakup Bär and Eiswirth (1993b). Similarly, the Puschino model is extensively utilized in numerical simulations of excitable media Panfilov and Rudenko (1987). The Karma model, another variant, was developed to demonstrate spiral wave breakup due to alternans Karma (1993). It generates an action potential with a more realistic fast upstroke and slow recovery compared to the standard FHN model. Additionally, the Aliev-Panfilov model adapts the FHN framework to capture the rate-dependent variations in action potential duration Aliev and Panfilov (1996).
@@ II.2 Linear stability of the stationary solutions
The position of these bifurcations vary as a function of the control parameters of the system $a,b$ and $\varepsilon$, leading to the emergence of different dynamical behaviors and regimes. An example of the distribution of these regimes and the bifurcations which define them is illustrated in the $(a,b)$-phase diagram of Fig. 5B for $\varepsilon=0.01$.
@@ III The diffusively coupled FitzHugh-Nagumo system
Originally conceived for neuroscience applications, the FHN model FitzHugh (1961) and its spatial extension Nagumo et al. (1962) have been used beyond their initial scope, finding relevance across various disciplines. In cardiology, the model is celebrated for its ability to simulate spiral wave dynamics Franzone et al. (2006); Davidenko et al. (1995); Courtemanche et al. (1990); Keener (1988); Keener and Sneyd (2009); Aliev and Panfilov (1996); Biktashev (2002); Pertsov et al. (1993); Nash and Panfilov (2004); Watanabe et al. (2004); Sermesant et al. (2006); Cherubini et al. (2008); Glass and Josephson (1995); Le and Sotiropoulos (2013); Panfilov and Keener (1995); Nagai et al. (2000), while more generally in biology, it is recognized for generating traveling waves that effectively convey information Gelens et al. (2014); Noguchi (2023); Nolet et al. (2020a); Gambino et al. (2022, 2023); Cartwright et al. (1999); Duffy et al. (1998); Sneyd et al. (2017b); Lee (1997); Gelens et al. (2015). The model’s utility extends to chemistry and computational sciences Moayeri et al. (2020); Dehghan and Narimani (2018); Fuselier and Wright (2013); Rogers and McCulloch (1994); Kuramoto et al. (1998); Lin et al. (2004), where it aids in understanding complex systems. Beyond its interdisciplinary applications, the FHN model has been instrumental in investigating mathematical properties Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Keener (1986a); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Ohta et al. (1989); Iqbal et al. (2017); Elphick et al. (1990); Meron (1992); Petrich and Goldstein (1994); Hagberg and Meron (1994a); Goldstein et al. (1996); Vasiev (2004); Nekhamkina and Sheintuch (2006); Panfilov and Hogeweg (1993); Winfree (1991a); Zheng and Shen (2015); Tsyganov et al. (2014); Rombouts and Gelens (2020); Winfree (1991b); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Zaritski et al. (2004); Bär and Eiswirth (1993a); Kuznetsov et al. (2017); Shepelev et al. (2017); Browne et al. (2008); Guckenheimer and Kuehn (2010, 2009); Perc (2005b); Shima and Kuramoto (2004); Hakim and Karma (1999); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Métens et al. (1997); Zemskov and Epstein (2010); Hastings (1982); Lin et al. (2004); Bode (1997); Tobias and Knobloch (1998); Marée and Panfilov (1997); Shoji et al. (2007); Yang et al. (2003), establishing it as a fundamental model in theoretical studies. This section aims to provide a theoretical exploration into the existence and linear stability of stationary homogeneous solutions and the emergence of spatially structured dynamics such as traveling waves and extended patterns. For in-depth mathematical treatments, readers are directed to Cross and Greenside (2009).
@@ III.6.2 Stationary patterns in two spatial dimensions
Stationary patterns, stemming from Turing bifurcations, have been a significant area of study. Various researchers have looked into the genesis of these spatial formations, formulating amplitude equations Iqbal et al. (2017); Zheng and Shen (2015). A notable challenge is the close proximity of Turing and Hopf bifurcations, complicating pattern identification. To circumvent this, strategies involve operating near the Turing threshold or employing white noise to steer the system towards stable pattern branches. Our numerical simulations reveal a diversity of stable spatial patterns across different control parameters, specifically for $b=1.26$, $\varepsilon=0.5$, $D_{u}=1$, and $D_{v}=5$, as illustrated in Fig. 13. These patterns, transitioning from negative to positive spots through labyrinthine structures, underscore the labyrinth configuration’s robust stability for the chosen parameters. Notably, Goldstein’s work on interface growth within the FHN framework Goldstein et al. (1996), inspired by chemical front interactions Lee et al. (1993), aligns with our findings in the labyrinth domains (Fig. 13B).
@@ III.6.5 Spiral waves
For the latter, we focus on spirals, which have been widely studied in the FHN model for mathematical purposes Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993a); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) and for their applicability in cardiology Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), where spiral behavior has been associated with arrhythmias. From the simple spirals studied by Erhaud and Winfree Meron (1992); Winfree (1991b) to the complex collisions studied by the latter Winfree (1991a), and through spiral turbulence emerging from the loss of stability of these structures Zaritski et al. (2004); Bär and Eiswirth (1993a), the FHN model has been utilized to characterize several regimes associated with spiral behavior. In Fig. 14B, we illustrate the creation of a stable spiral arising from a phase defect induced by considering horizontal gradient-like initial conditions for $u$ and horizontal gradient-like initial conditions for $v$ in an oscillatory regime ($a=0.1$, $b=1$, $\varepsilon=0.005$, $D_{u}=1$, $D_{v}=1$).
Spiral wave patterns, particularly relevant in cardiological contexts due to their association with arrhythmic behaviors Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), represent another facet of the FHN model’s versatility. From the foundational works on simple spirals Meron (1992); Winfree (1991b) to the studies on spiral turbulence Zaritski et al. (2004); Bär and Eiswirth (1993a) and collision dynamics Winfree (1991a), the model serves as a vital tool in understanding the nuances of spiral behavior. Fig. 14B demonstrates the formation of a stable spiral, a phenomenon induced by specific initial conditions (here a phase defect) within an oscillatory setting. This insight into spiral formation within the FHN model not only advances mathematical explorations Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993a); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) but also has profound implications in the study of cardiac arrhythmias Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), offering potential avenues for therapeutic interventions.
@@ IV Discretely coupled FitzHugh-Nagumo equations
F Goetze and Lai (2022); Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Stefanescu and Jirsa (2008); Li et al. (2021); Binczak et al. (2006); Yanagita et al. (2005) F T Wang et al. (2009); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Toral et al. (2003b); Yu et al. (2021); Brandstetter et al. (2010); Panchuk et al. (2013); Ciszak et al. (2003); Schöll et al. (2009); Saha and Feudel (2017) T F Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner (2004); Brown et al. (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005a); Takembo et al. (2019) F Lei and Yang (2021); Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007); Steur et al. (2009); Zhang and Liao (2017); Chouzouris et al. (2018); Zaks et al. (2005); Vodeneev et al. (2011); Sukhov et al. (2011); Schmidt et al. (2017); García-Ojalvo and Schimansky-Geier (1999); Cartwright (2000); Gerster et al. (2020); Scialla et al. (2021); Ullner et al. (2003b) T Burić and Todorović (2003); Kori et al. (2012); Perlikowski et al. (2010) T Sawicki et al. (2019); Schülen et al. (2019) Discretely coupled ODEs, describing various network configurations, have garnered substantial attention in diverse fields such as neuroscience Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020), electrical systems Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006); Takembo et al. (2019); Ciszak et al. (2003), and biology Kori et al. (2012); Vodeneev et al. (2011); Sukhov et al. (2011); Cartwright (2000), owing to their relevance and the broad spectrum of topologies and coupling terms they accommodate. This versatility allows for the exploration of various phenomena, enriching the mathematical study of networks Vaidyanathan (2015); Hoff et al. (2014); Lei and Yang (2021); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Lindner (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005a); Steur et al. (2009); Zaks et al. (2005); Schmidt et al. (2017); García-Ojalvo and Schimansky-Geier (1999); Gerster et al. (2020); Yanagita et al. (2005).
@@ V Conclusions
We structured our analysis into three primary sections. Firstly, we examined the original FHN model [Eq. (8)], discussing widely observed dynamical regimes such as monostability, multistability, relaxation oscillations, and excitability. We examined the role of local and global bifurcations in shaping these regimes, emphasizing the importance of time scale separation. Secondly, we explored the diffusively coupled FHN model [Eq. (10)], introducing spatial coupling through diffusion. Through theoretical analysis, we investigated stationary homogeneous solutions, their linear stability, and spatially structured dynamic solutions, including traveling structures and spatially extended patterns. We studied the emergence of a Turing instability and the resulting spatially structured Turing patterns. Additionally, we examined front solutions, localized states, traveling pulses, and pacemaker-driven waves within the oscillatory domain, highlighting the richness of patterns that arose in different spatial dimensions. Lastly, we explored discretely coupled FHN equations [Eq. (11)]. This is the broadest category as here one can consider a multitude of different network topologies and coupling terms. We focussed ourselves on synchronization properties in two coupled FHN modules, the existence of traveling waves when transitioning from continuous diffusive coupling to discrete coupling, and the emergence of chimera states characterized by spatiotemporal patterns of coherent and incoherent behavior.